# GrEmLIn: A Repository of Green Baseline Embeddings for 87 Low-Resource Languages Injected with Multilingual Graph Knowledge

### v2

@@ Abstract
Contextualized embeddings based on large language models (LLMs) are available for various languages, but their coverage is often limited for lower resourced languages. Using LLMs for such languages is often difficult due to a high computational cost; not only during training, but also during inference. Static word embeddings are much more resource-efficient ("green"), and thus still provide value, particularly for very low-resource languages. There is, however, a notable lack of comprehensive repositories with such embeddings for diverse languages. To address this gap, we present GrEmLIn, a centralized repository of green, static baseline embeddings for 87 mid- and low-resource languages. We compute GrEmLIn embeddings with a novel method that enhances GloVe embeddings by integrating multilingual graph knowledge, which makes our static embeddings competitive with LLM representations, while being parameter-free at inference time. Our experiments demonstrate that GrEmLIn embeddings outperform state-of-the-art contextualized embeddings from E5 on the task of lexical similarity. They remain competitive in extrinsic evaluation tasks like sentiment analysis and natural language inference, with average performance gaps of just 5-10% or less compared to state-of-the-art models, given a sufficient vocabulary overlap with the target task, and underperform only on topic classification. Our code and embeddings are publicly available at https://huggingface.co/DFKI.
@@
Daniil Gurgurov  Rishu Kumar  Simon Ostermann German Research Center for Artificial Intelligence (DFKI) {daniil.gurgurov, rishu.kumar, simon.ostermann}@dfki.de
@@ 1 Introduction
While contextual representations like the ones based on BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and GPT Radford et al. (2019) nowadays provide better performance than static embeddings in many tasks, their training is computationally expensive Strubell et al. (2019); Bommasani et al. (2021) and ineffective for data-scarce languages due to their data hunger and the curse of multilinguality Conneau et al. (2020). Also, they still require hardware during runtime, as embeddings need to be computed based on a forward pass for each new text that is processed. This is often prohibitive in low-resource (hardware) scenarios, and inefficient in terms of energy use.
@@
As for large language models (LLMs), the training of word embeddings suffers from the lack of high-quality data in low-resource languages (to a smaller degree). Incorporating other types of data for improving word representations is thus beneficial especially for low-resource languages. Knowledge graphs provide such an alternative to textual knowledge, with rich semantic and multilingual sources of information, including synonyms, antonyms, morphological forms, definitions, etimological relations, translations, and more Miller (1995); Navigli and Ponzetto (2012); Speer et al. (2017). Such structured and cross-lingual information can be used to improve the quality of classical word representations Faruqui et al. (2014); Sakketou and Ampazis (2020), which are only trained on co-occurence statistics.
@@
In summary, our contributions in this work are two-fold: First, we present GrEmLIn, a centralized resource of static word embeddings for 87 mid- and low-resource languages, specifically focusing on word embeddings trained with GloVe Pennington et al. (2014). Second, we propose an effective method to improve embeddings by incorporating more knowledge in the form of multilingual knowledge graphs, which is especially important for low-resource languages, where resources are usually very scarce [^1] https://github.com/d-gurgurov/GrEmLIn-Green-Embeddings-LRLs.
@@ 2 Related Work
There are various methods to improve word embeddings by incorporating external knowledge graphs or semantic networks Dieudonat et al. (2020). Retrofitting Faruqui et al. (2014) is a post-processing technique that adjusts pre-trained word embeddings using information from knowledge graphs or semantic lexicons. The key idea is to infer new vectors that are close to their original embeddings while also being close to their neighbors in the graph or lexicon. This is achieved by minimizing an objective function that balances the distance between the new vectors and the original embeddings, as well as the distance between connected nodes. Expanded retrofitting Speer et al. (2017), used for ConceptNet Numberbatch, optimizes over a larger vocabulary including terms from the knowledge graph not present in the original embeddings, but it still does not retrofit all the words in the original embedding space. Other existing methods that integrate contextualized embeddings with knowledge graph embeddings often use attention mechanisms, as demonstrated by works such as Peters et al. (2019) and Zhang et al. (2019). These methods specifically enhance BERT embeddings by incorporating external knowledge bases.
@@ 3 Method
We propose a method for merging GloVe embeddings with graph-based embeddings derived from ConceptNet knowledge, while preserving the vocabulary size of GloVe, following two steps: First, we use Singular Value Decomposition (SVD) on concatenated word embeddings from GloVe and PPMI-based graph embeddings to generate a shared embedding space. We do so for the part of the vocabulary that is shared between GloVe and the knowledge graph. Second, we learn a linear transformation from GloVe into this joined space to obtain embeddings for all words in the original GloVe vocabulary.
@@ 3.1 GloVe Embeddings
We train GloVe embeddings using the original code. The model is trained by stochastically sampling nonzero elements from the co-occurrence matrix over 100 iterations, to produce 300-dimensional vectors. We use a context window of 10 words to the left and 10 words to the right. Words with fewer than 5 co-occurrences are excluded for languages with over 1 million tokens in the training data, and the threshold is set to 2 for languages with smaller datasets. We use data from CC100222https://huggingface.co/datasets/cc100 Wenzek et al. (2020); Conneau et al. (2020) for training the static word embeddings. We set $x_{max}=100$, $\alpha=\frac{3}{4}$, and use AdaGrad optimization Duchi et al. (2011) with an initial learning rate of 0.05.
@@ 4.2 Evaluation Data
Lastly, we evaluate the embeddings on the NLI task, using the XNLI dataset Conneau et al. (2018). The XNLI dataset provides multilingual NLI examples for 15 languages, and for our experiments, we selected 4 of these languages: Swahili, Urdu, Greek, Thai, and Bulgarian. These languages were selected based on the availability of our GloVe, PPMI-enhanced embeddings, and the NLI dataset. Evaluating embeddings on the NLI task tests their ability to understand logical relationships between sentence pairs, an important capability for higher-level NLU tasks. Due to the simplicity of our models, we only utilize validation and test splits, consisting of 2,490 and 5,010 examples, respectively, for training and testing, excluding the original training split of nearly 400,000 examples.
@@ Acknowledgments
The work was supported by DisAI – Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164.

### v3

@@ Abstract
Contextualized embeddings based on large language models (LLMs) are available for various languages, but their coverage is often limited for lower resourced languages. Using LLMs for such languages is often difficult due to a high computational cost; not only during training, but also during inference. Static word embeddings are much more resource-efficient ("green"), and thus still provide value, particularly for very low-resource languages. There is, however, a notable lack of comprehensive repositories with such embeddings for diverse languages. To address this gap, we present GrEmLIn, a centralized repository of green, static baseline embeddings for 87 mid- and low-resource languages. We compute GrEmLIn embeddings with a novel method that enhances GloVe embeddings by integrating multilingual graph knowledge, which makes our static embeddings competitive with LLM representations, while being parameter-free at inference time. Our experiments demonstrate that GrEmLIn embeddings outperform state-of-the-art contextualized embeddings from E5 on the task of lexical similarity. They remain competitive in extrinsic evaluation tasks like sentiment analysis and natural language inference, with average performance gaps of just 5-10% or less compared to state-of-the-art models, given a sufficient vocabulary overlap with the target task, and underperform only on topic classification. Our code and embeddings are publicly available at https://huggingface.co/DFKI [^1] All vectors are available on Huggingface as single model pages. Each page starts with DFKI/glove..
@@
Daniil Gurgurov1  Rishu Kumar1  Simon Ostermann1,2 1German Research Center for Artificial Intelligence (DFKI) 2Centre for European Research in Trusted AI (CERTAIN) {daniil.gurgurov, rishu.kumar, simon.ostermann}@dfki.de
@@ 1 Introduction
While contextual representations like the ones based on BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and GPT Radford et al. (2019) nowadays provide better performance than static embeddings in many tasks, their training is computationally expensive Strubell et al. (2019); Bommasani et al. (2021) and ineffective for data-scarce languages due to their data hunger and the curse of multilinguality Conneau et al. (2020). Some approaches of efficient adaptation of especially large language models (LLMs) to languages other than English have been investigated in recent years Pfeiffer et al. (2020); Vykopal et al. (2024). However, even such approaches still require hardware during runtime, as embeddings need to be computed based on a forward pass for each new text that is processed. This is often prohibitive in low-resource (hardware) scenarios, and inefficient in terms of energy use. Also, such approaches are often not tailored to low-resource languages.
@@
As for LLMs, the training of word embeddings suffers from the lack of high-quality data in low-resource languages (to a smaller degree). Incorporating other types of data for improving word representations is thus beneficial especially for low-resource languages. Knowledge graphs provide such an alternative to textual knowledge, with rich semantic and multilingual sources of information, including synonyms, antonyms, morphological forms, definitions, etimological relations, translations, and more Miller (1995); Navigli and Ponzetto (2012); Speer et al. (2017). Such structured and cross-lingual information can be used to improve the quality of classical word representations Faruqui et al. (2014); Sakketou and Ampazis (2020), which are only trained on co-occurence statistics.
@@
In summary, our contributions in this work are two-fold: First, we present GrEmLIn, a centralized resource of static word embeddings for 87 mid- and low-resource languages, specifically focusing on word embeddings trained with GloVe Pennington et al. (2014). Second, we propose an effective method to improve embeddings by incorporating more knowledge in the form of multilingual knowledge graphs, which is especially important for low-resource languages, where resources are usually very scarce. Our code is publicly available on GitHub [^2] https://github.com/d-gurgurov/GrEmLIn-Green-Embeddings-LRLs.
@@ 2 Related Work
There are various methods to improve word embeddings by incorporating external knowledge graphs or semantic networks Dieudonat et al. (2020). Retrofitting Faruqui et al. (2014) is a post-processing technique that adjusts pre-trained word embeddings using information from knowledge graphs or semantic lexicons. The key idea is to infer new vectors that are close to their original embeddings while also being close to their neighbors in the graph or lexicon. This is achieved by minimizing an objective function that balances the distance between the new vectors and the original embeddings, as well as the distance between connected nodes. Expanded retrofitting Speer et al. (2017), used for ConceptNet Numberbatch, optimizes over a larger vocabulary including terms from the knowledge graph not present in the original embeddings, but it still does not retrofit all the words in the original embedding space. Other existing methods that integrate contextualized embeddings with knowledge graph embeddings often use attention mechanisms, as demonstrated by works such as Peters et al. (2019), Zhang et al. (2019), and Gurgurov et al. (2024). These methods specifically enhance BERT embeddings by incorporating external knowledge bases.
@@ 3 Method
We propose a method for merging GloVe embeddings with graph-based embeddings derived from ConceptNet knowledge, while preserving the vocabulary size of GloVe, following two steps: First, we use singular value decomposition (SVD) Eckart and Young (1936) on concatenated word embeddings from GloVe and pointwise mutual information (PMI) based graph embeddings Speer et al. (2017) to generate a shared embedding space. We do so for the part of the vocabulary that is shared between GloVe and the knowledge graph. Second, we learn a linear transformation from GloVe into this joined space to obtain embeddings for all words in the original GloVe vocabulary.
@@ 3.1 GloVe Embeddings
We train GloVe embeddings using the original code. The model is trained by stochastically sampling nonzero elements from the co-occurrence matrix over 100 iterations, to produce 300-dimensional vectors. We use a context window of 10 words to the left and 10 words to the right. Words with fewer than 5 co-occurrences are excluded for languages with over 1 million tokens in the training data, and the threshold is set to 2 for languages with smaller datasets. We use data from CC100333https://huggingface.co/datasets/cc100 Wenzek et al. (2020); Conneau et al. (2020) for training the static word embeddings. We set $x_{max}=100$, $\alpha=\frac{3}{4}$, and use AdaGrad optimization Duchi et al. (2011) with an initial learning rate of 0.05.
@@ 3.4 Linear Transformation
[…]
$W$ can be learned by solving the following optimization problem:
$\min_{W}\sum_{i=1}^{n}\|Wx_{i}-z_{i}\|^{2}$
which we solve as a linear regression problem with stochastic gradient descent optimization.
@@ 4.2 Evaluation Data
Lastly, we evaluate the embeddings on the NLI task, using the XNLI dataset Conneau et al. (2018). The XNLI dataset provides multilingual NLI examples for 15 languages, and for our experiments, we selected 5 of these languages: Swahili, Urdu, Greek, Thai, and Bulgarian. These languages were selected based on the availability of our GloVe, PPMI-enhanced embeddings, and the NLI dataset. Evaluating embeddings on the NLI task tests their ability to understand logical relationships between sentence pairs, an important capability for higher-level NLU tasks. Due to the simplicity of our models, we only utilize validation and test splits, consisting of 2,490 and 5,010 examples, respectively, for training and testing, excluding the original training split of nearly 400,000 examples.
@@ Acknowledgments
We thank the anonymous reviewers for their helpful feedback. This work was supported by DisAI – Improving scientific excellence and creativity in combating disinformation with artificial intelligence and language technologies, a Horizon Europe-funded project under GA No. 101079164, and by the German Ministry of Education and Research (BMBF) as part of the project TRAILS (01IW24005).